A football is kicked vertically upward from a height of 4 feet with an initial speed of 75 ft per second. What is the height of the ball 2 seconds after it was kicked from the ground.
H=4+75t-16t^2

Answer by rapaljer(4671)

H=4+75t-16t^2, where t=2 sec
H=4+75*2-16*2^2
H=4+ 150 - 16*4
H=4+150-64
H=154-64=90 feet

Or you can do it with a graphing calculator if you want to. If anyone is interested in this approach, let me know. Send me an Email at rapaljer@scc-fl.edu.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus